OpenStudy (anonymous):

Please help!!!!! What is the answer to: (x-5)/(3x+4) - (3x-5)/(3x+4) Thanks!

OpenStudy (anonymous):

denominators are the same, subtract in the numerator, use parentheses

OpenStudy (anonymous):

\[\frac{x-5}{3x+4}-\frac{3x-5}{3x+4}=\frac{x-5-(3x-5)}{3x+4}\]

OpenStudy (anonymous):

no remove the parentheses and combine like terms in the numerator \[\frac{x-5-(3x-5)}{3x+4}=\frac{x-5-3x+5}{3x+4}=\frac{-2x}{3x+4}\]

OpenStudy (anonymous):

Wow! Thanks so much!

OpenStudy (anonymous):

yw